OpenStudy (anonymous):

What are the solutions to 6x^2 = 18x?

OpenStudy (anonymous):

6x^2-18x=0 6x(x-3)=0 Therefore, x=0 or x=3

OpenStudy (anonymous):

Thanks! I'm going over questions I got wrong on a test.

OpenStudy (anonymous):

It's okay. What class are you in, btw?

OpenStudy (anonymous):

Algebra 1

OpenStudy (anonymous):

I mean grade.

OpenStudy (anonymous):

8th

OpenStudy (anonymous):

6x^2-18x=0 6x(x-3)=0 6x^2 - 18x = 0 6x^2 - 18x = 0 6x^2 = 18x x^2 = 3x x^2/x = 3x/x x = 3